OpenStudy (anonymous):

Determine whether or not the triangle with vertices A(3, 5), B(1, 4), and C(6, 2) is a right triangle.

hero (hero):

Have you plotted the points yet?

OpenStudy (anonymous):

I made a rough sketch

hero (hero):

You need a precise sketch using graph paper. And when you create those points, draw lines connecting them to form a triangle. Then count the length of the sides and test to make sure pythagorean theorem works. When you do that, you will have shown whether or not the vertices form a right triangle.

OpenStudy (anonymous):

It HAS to be drawn?

OpenStudy (mimi_x3):

lol, just use the distance formula

OpenStudy (anonymous):

:P Whats the distance formula?

hero (hero):

So to summarize, the steps are to 1. Plot the points 2. Connect the points to form a triangle 3. Use pythagorean theorem to verify the right triangle. 4. Doesn't HAVE to be drawn to do this.

hero (hero):

But it is easier when you draw

OpenStudy (anonymous):

Yea, from my drawing, I can deduce that it does not look like a right angle triangle :P

OpenStudy (mimi_x3):

drawing cannot be precise you have to use graph paper; just use the distance formula

OpenStudy (anonymous):

WHATS THE DISTANCE FORMULA?

hero (hero):

The distance formula is just a variation of the pythagorean theorem

OpenStudy (anonymous):

Hmm... So how can I answer this question without drawing because I dont have any graph paper

hero (hero):

Instead of a^2 + b^2 = c^2, you have a^2 + b^2 = d^2 and you solve for d

OpenStudy (mimi_x3):

\[\large d= \sqrt{(x_{2}-x_{1})^{2}+(y_{2}-y_{1})^{2}} \]

hero (hero):

except for a = (x_2 -x_1) and b = (y_2 - y_1)

OpenStudy (anonymous):

not necessarily just find the slopes if two are perp then you're good togo

hero (hero):

Outkast ignores the part where it says "right triangle".

hero (hero):

right triangle implies the use of pythagorean theorem

hero (hero):

you can prove two slopes are perpendicular but that's only part of it.

OpenStudy (anonymous):